Pasqually Mineral Water, Inc., will pay a quarterly dividend per share of $0.25 at the end of each of the next 16 quarters. Thereafter, the dividend will grow at aquarterly rate of 1.5 percent, forever. The appropriate rate of return on the stock is 12 percent, compounded quarterly.What is the current stock price
Answer #1


Pasqually Mineral Water, Inc., will pay a quarterly dividend per share of $0.25 at the end of each of the next16 quarters. Thereafter, the dividend will grow at a quarterly rate of 1.5 percent,forever. The appropriate rate of return on the stock is 12 percent, compounded quarterly.

What is the current stock price = 0.388 DOLLARS -ANSWER

= 0.25 /(12% -1.5% ) =2.38

current stock price =2.38 X 1.12 ^16 =0.388


answered by: saffy
Answer #2
value after 16 qtr= div per share/(rate of return - dividend grwth rate) =0.25/(.12-.015)=$2.38

current value=2.38/1.12^16=2.38/6.13=$.388 ans
answered by: kerrae-rae
Answer #3
current price = 0.025/(0.12+0.015)=0.185 $
ans is $0.185
answered by: kadian